OpenStudy (anonymous):

Y=Xy^4+x^2y=x+3y can some help me derive this

OpenStudy (anonymous):

Don't understand. What is the question?

myininaya (myininaya):

do you find the derivative? derive does not equal derivative. do you mean to have one equality?

OpenStudy (anonymous):

i thinks am so suppose to solve using implicit

OpenStudy (anonymous):

that what i dont understad either the book just says solve and am doing derivatives

OpenStudy (anonymous):

Implicit means the y is embedded so when you write it, you don't write y=.... Just writ3 xy^4+x^2y=x=3y

OpenStudy (anonymous):

this is what the book is giving me as an answer if it helps 1-y^4-2xy/4x^3+x^2-3

OpenStudy (anonymous):

In implicit differentiation you differentiate each item (Some will be product rule) Where ever you find derivative of y (including in product rule you write dy/dx. For ex derivative of 2y is 2dy/dx. Then you solve equation for dy/dx

OpenStudy (anonymous):

so x+3y would be 3 + x

OpenStudy (anonymous):

It is usually in an equation. derivative of x + 3y=0 would be 1 + 3 dy/dx=0. (Not to be confused with partial differentiation where you differentiate only one and don't differentiate the other.)

OpenStudy (anonymous):

To finish up 3dy/dx=-1 dy/dx =-1/3